# Mixed Warm Dark Matter Constraints using Milky Way Satellite Galaxy Counts

### v2

@@ Acknowledgements.
The authors would like to thank Ethan Nadler, Rui An, Vera Gluscevic, and Andrew Benson for the helpful discussions about MWDM scenarios and for the help with implementing the galaxy–halo connection model code. They would also like to thank Jacopo Ghiglieri for his comments about the different sterile neutrino codes, and Daniel Gilman, Chihway Chang, and Ava Polzin for extended discussion on the statistical formalism used to derive constraints. This work was partially supported by the National Science Foundation (NSF) through AST-2307126. AD is supported by the Kavli Institute for Cosmological physics at the University of Chicago through an endowment from the Kavli Foundation and its founder Fred Kavli. This manuscript has been authored by Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics.

### v3

@@ Acknowledgements.
The author would like to thank the anonymous referee for the many useful comments that helped us improve this manuscript. They would also like to thank Ethan Nadler, Rui An, Vera Gluscevic, and Andrew Benson for the helpful discussions about MWDM scenarios, Jacopo Ghiglieri for his comments about the different sterile neutrino codes, and Daniel Gilman, Chihway Chang, and Ava Polzin for extended discussion on the statistical formalism used to derive constraints. This work was partially supported by the National Science Foundation (NSF) through AST-2307126. AD is supported by the Kavli Institute for Cosmological physics at the University of Chicago through an endowment from the Kavli Foundation and its founder Fred Kavli. This manuscript has been authored by Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics.